OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!! PLEASE PLEASEE PLEASE *WITH PUPPY EYES* I WILL GIVE MEDAL AND BE FAN

OpenStudy (anonymous):

what is your question

OpenStudy (anonymous):

square root of -81/100

OpenStudy (anonymous):

square root(-81) / 100 = 0.09 i

OpenStudy (anonymous):

as a fraction plz

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

2 questions after this one btw

OpenStudy (anonymous):

go on this and you will get your answer www.youtube.com/watch?v=Thxzb35HQlw‎

OpenStudy (shamil98):

\[\huge \frac{ \sqrt{-81} }{\sqrt{100} } = \frac{ 9i }{ 10 }\]

OpenStudy (anonymous):

THX SHAMIL

OpenStudy (anonymous):

sorry if i did not help

OpenStudy (anonymous):

u did plus two more questions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

What are the solutions to the equation x² = 256? Choose all answers that are correct. A. 128 B. –128 C. 16 D. –16

OpenStudy (anonymous):

D

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

well bye

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

what :D

OpenStudy (anonymous):

there's multiple choice and plus i need 1 more question

OpenStudy (anonymous):

ok hit it :D

OpenStudy (anonymous):

what other answer do u think is it

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

thx

jimthompson5910 (jim_thompson5910):

x^2 = 256 x = +-sqrt(256) ... don't forget the plus/minus x = sqrt(256) or x = -sqrt(256) x = 16 or x = -16

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

What is ur question :D

OpenStudy (anonymous):

so.......................what is your question

OpenStudy (anonymous):

oh sorry What are the solutions to the equation y² –5 = 31? Choose all answers that are correct. A. 18 B. 6 C. –18 D. –6

OpenStudy (anonymous):

-18

jimthompson5910 (jim_thompson5910):

hint: y^2 - 5 = 31 turns into y^2 = 36 when you add 5 to both sides

OpenStudy (anonymous):

forget what i said he is right^

OpenStudy (anonymous):

ok thx anyway angel

OpenStudy (anonymous):

ok gtg bye

OpenStudy (anonymous):

bye :D

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

what :D

OpenStudy (anonymous):

so what answer choices is that

OpenStudy (anonymous):

for what

OpenStudy (anonymous):

i am confused what do you mean

OpenStudy (anonymous):

what is the anwser

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thx :D